A block of mass M is resting on a smooth horizontal surface on one end of a uniform rope of mass M by 368 to the block which is pulled in horizontal direction applying force f at the other at the tension in the middle of the Rope is?

Grade:11

2 Answers

Harsh
4 years ago
Please write the question in proper manner.
I am unable to understand
Wht is 368 in the question.
Please re write the question i will post ans as soon as possible
Vikas TU
4 years ago
Dear student 
I assumed the mass is m/3 , an solved the question. 
the total mass of the system becomes m/3 + m = 4m/3
therefore acceleration  of the system = f/m=  3f/4m
let the length of the rope be l so,
for l mass is m/3 + m and for 
l/2 mass is m/6 + m    that bis 7m/6
 
therefor tension in the mid point =m *a
=7m/6*3f/4m
= 21f/24